OpenStudy (anonymous):

What is the solution to the equation 4 –2x ≈ 3 ?

OpenStudy (smarty):

-2x=-1 x=1/2

OpenStudy (anonymous):

What does the symbol after -2x means???

OpenStudy (smarty):

= sign.

OpenStudy (anonymous):

You take 4 the other side of the equal sign. So it will be 3-4 which is equal to -1. And then to remove -2 from x, you divide -2 both sides which will give you the final value of x as 1/2 or 0.5

OpenStudy (jiteshmeghwal9):

4-2x=3 =>-2x=3-4 =>-2x=-1 =>x=-1/-2=1/2.